# A Unified, Practical, and Understandable Summary of Non-transactional Consistency Levels in Distributed Replication

### v2

@@ Abstract.
We present a summary of practical non-transactional consistency levels in the context of distributed data replication. Unlike prior work, we build upon a simple Shared Object Pool (SOP) model and define consistency levels in a unified framework centered around the concept of ordering. This model naturally reflects modern cloud object storage services and is thus easy to understand. We show that a consistency level can be intuitively defined by specifying two types of constraints on the validity of orderings allowed by the level: convergence, which bounds the lineage shape of the ordering, and relationship, which bounds the relative positions between operations. We give examples of representative protocols and systems, and discuss their availability upper bound.
@@ 1. Introduction
Section 2 describes our problem model setup, defines logical ordering, and explains the meaning of non-transactional consistency within this context. Section 3 defines all variants of ordering validity constraints. Section 4 presents the hierarchy of selected consistency levels, dissects their ordering validity guarantees, explains their practical differences, and gives examples of representative protocols and systems. Section 5 discusses the availability upper bound of each level in the presence of network partitioning. Section 6 concludes this paper.
@@ 2.1. Shared Object Pool (SOP) Model
We consider a storage service shared by multiple clients, as shown in Figure 1. The service appears as a pool of objects. Each object has a unique name and contains a value. The only way to learn about an object’s value is through the result of a client read operation, which we introduce below. Objects are not necessarily stored as physical bytes on physical machines; the SOP model is entirely conceptual and is agnostic to any actual design of protocols and implementation of systems.
Clients are single-threaded, closed-loop entities that invoke operations on the service. When a client $c$ issues an operation $p$, it will block until the acknowledgment of $p$ by the service. Multi-threaded or asynchronous client implementations should be modeled as multiple SOP clients.
An operation is of one of the following three types:
@@
The service maintains a possibly partial ordering $O$ of all operations that have been acknowledged. The ordering $O$ captures all dependencies between operations enforced by the service and therefore materializes the result of each operation. Given a workload of operations generated by clients, whether an ordering is acceptable or not is decided by its validity constraints. Modeling the ordering validity constraints guaranteed by the service effectively models its interface semantics, hence its consistency level. The following three subsections explain the meaning of workload, ordering, and consistency, respectively.
@@ 2.4. Meaning of Consistency
The consistency level of the storage service is determined by which orderings of operations are considered valid given any physical timeline workload. In other words, the consistency level enforces what validity constraints must be held on the ordering given any workload. A stronger consistency level imposes more constraints than a weaker one and therefore disallows more orderings, exposing an interface that is more restrictive in the protocol design space and in the meantime easier to use by clients. In contrast, a weaker consistency level relaxes certain constraints and opens up new opportunities in the protocol design space, however providing weaker semantic guarantees for clients.
An ordering represents logical dependencies among operations, similar to Lamport’s definition of logical clock on events (Lamport, 1978b), and does not necessarily capture physical time. In fact, whether physical time is respected or not is one of the validity constraints that differentiate several consistency levels, as we demonstrate in Section 4. Our SOP model shares similarities with the specification framework for replicated data types proposed by Burckhardt et al. (Gotsman et al., 2014); the differences are that we simplify the notion of ordering and cover stronger consistency levels (rather than focusing only on causal and eventual consistency models).
@@ 3.1.2. Convergent Partial Order (CPO)
In addition, the result of a read must be strongly convergent (Viotti and Vukolić, 2016), meaning that it must observe all operations to the same object that immediately precede it. If multiple operations with different values to the same object all immediately precede the read and they are unordered with each other, then the read must return the set of all these values (or a reduced value over the set by applying a convergent reduction function, as described in Section 2.1).
Below is an example ordering that satisfies CPO (but does not satisfy SO):
@@ 3.1.3. Non-convergent Partial Order (NPO)
Below is an example ordering that satisfies NPO (but does not satisfy CPO):
@@ 3.2.2. Causal (CASL)
For instance, the following is an SO ordering that satisfies CASL (but does not satisfy RT), given the same example timeline presented in Section 3.2.1:
@@
Session Guarantees. A popular approach to interpreting causality, as first described in (Terry et al., 1994), is to think from each client’s perspective (termed a session) and decompose the CASL constraint into four session guarantees:
@@ 3.2.3. First-In-First-Out (FIFO)
For instance, the following is an SO ordering that satisfies FIFO (but does not satisfy CASL), given the same example timeline presented in Section 3.2.1:
@@ 3.2.4. None Relationship
An ordering could of course place no restrictions on the relative positions of operations. In this case, operations issued by the same client may get arbitrarily reordered. Writes by the same client may be visible to another client in a different order than issued, and a client’s read may fail to observe its own preceding write.
@@ 4.1. Linearizability
4.2. Sequential Consistency
For example, given the following physical timeline:
$c:$$d:$$c\text{W}x{\angle}1$$d\text{W}x{\angle}2$$d\text{R}x$
A linearizable ordering must be SO and RT:
$|c\text{W}x{\angle}1|$ $|d\text{W}x{\angle}2|$ $|d\text{R}x{:}2|$
$|d\text{W}x{\angle}2|$ $|c\text{W}x{\angle}1|$ $|d\text{R}x{:}1|$
@@ 4.2. Sequential Consistency
In the example depicted by Figure 3, it could be that client $c$ first issues a write of value 1 to object $x$ and waits for its acknowledgment. It then sends a message to client $d$ through an external inter-client channel saying “I have finished my write to $x$ and you can go ahead to operate on $x$.” Client $d$ then issues its own write of value 2 and expects to read out 2 afterwards. However, since the message from $c$ to $d$ is external to the service, a sequentially consistent service may reorder $d$’s write ahead of $c$’s, and return value 1 for $d$’s read.
@@
A service that provides linearizability will be able to capture such implicit external dependencies because of the real-time property, as $|d\text{W}x{\angle}2|$ starts after $|c\text{W}x{\angle}1|$’s acknowledgment in physical time [^6] Note that this is not to be confused with the external consistency property in distributed transaction processing systems (Gifford, 1981; Corbett et al., 2013), which means that transactions are serialized into the same order as their commit order..
@@
Notice that given the result of $d$ reading 1 out of $x$ and $c$ reading 2 out of $y$, it is impossible to resolve an SO and CASL ordering across all six operations. This implies that a protocol that guarantees sequential consistency on each object may fail to come up with a global sequence of operations. In fact, such a protocol provides per-key sequential consistency (see Section 4.5.6).
@@
Adding support for read-modify-write (RMW) operations to such protocols is a non-trivial task (Burke et al., 2020). In particular, we cannot simply treat RMW operations in the same way as pure writes, because RMWs require a stable base value to determine the result of the read. Systems that demand compare-and-swap (CAS) operations (such as the LogOnce operation on shared logs (Guo et al., 2022)) may have to opt for a service that provides linearizability (or regular sequential consistency (Helt et al., 2021) as discussed in Section 4.5.1).
@@ 4.4. Eventual Consistency
Eventual consistency, as the name suggests, is a consistency level that only requires reads on an object to return a consistent value if no updates are being made to the object (Vogels, 2008). There is no relationship constraint between operations, meaning that any pair of operations issued by the same client are allowed to get reordered, let alone preserving causality, in the final ordering. Eventual consistency is widely adopted in geo-scale systems where the demand for high performance, scalability, and availability outweighs the need for timely consistency.
Eventual Convergence. Although eventual consistency is sometimes used interchangeably with weak consistency, it does impose one extra requirement on the service: the decided ordering must be convergent. In other words, after all the writers on an object become inactive and after all the writes become visible to readers, reads on the object must all return the same value. In our model, this is captured by the CPO constraint.
@@
Quiescent Consistency. A closely related, vaguely defined term is quiescent consistency (Herlihy and Shavit, 2011). In a commonly accepted definition, special periods of physical time are identified, during which no write operations are happening. Every such contiguous time period is called a quiescence period; all operations acknowledged ahead of the period are ordered before operations that start after the period. With this definition, quiescent consistency is weaker than eventual consistency, because it effectively makes no guarantees at all if a system-wide quiescence period never appears (Viotti and Vukolić, 2016).
Protocols & Systems. Eventual consistency is widely adopted by web-scale systems in the form of gossiping protocols and anti-entropy propagation (Schroeder et al., 1984; DeCandia et al., 2007; Elhemali et al., 2022; Lakshman and Malik, 2010). These systems value performance and scalability greatly and can tolerate inconsistencies.
@@ 4.5.1. Regular Sequential Consistency
Invariant-equivalence to Linearizability. It is shown that regular sequential consistency is invariant-equivalent to linearizability (Helt et al., 2021), meaning that: 1) it is local (see Section 4.2) and 2) it inherently supports RMW operations thanks to stable ordering of writes. However, it does not guarantee to capture external causality dependencies, making it still slightly weaker than linearizability. If external causality is not an issue, a linearizable replication system can seamlessly adopt regular sequential consistency to improve the performance of read-only operations.
@@ 4.5.2. Real-time Causal Consistency
Assuming that the system is composed of a set of symmetric message-passing replica nodes, Mahajan et al. have proven in (Mahajan et al., 2012) that real-time causal consistency is the strongest possible level that is achievable in an always-available, one-way convergent system (which is implied by our definition of sticky available in Section 5).
Fork-based Consistency Models. A family of fork-based consistency models has been developed to deal with Byzantine faults in a system containing untrusted replica nodes. For example, a fork-linearizable system ensures that if any two replicas have observed different orderings (i.e., forked by an adversary, even for one operation), then their writes will never be visible to each other afterwards (i.e., they cannot be joined again). Fork causal consistency is a family of consistency levels that weaken causal consistency to tolerate Byzantine replicas and enforce causal consistency among correct replicas (Mahajan et al., 2011).
@@ 4.5.3. Causal Consistency
Section 4.3 has explained causal and causal+ consistency. To recap, a service that provides causal consistency must give an ordering that is NPO and CASL given any physical timeline workload. Such an ordering captures all the potential causality dependencies between operations, but does not demand convergent conflict resolution, meaning that different clients are allowed to forever retrieve different values from reads on the same object.
As mentioned in Section 3.2.2, causal consistency can be defined exactly as the combination of the four session guarantees (Brzezinski et al., 2004; JEPSEN, 2016).
@@ 4.5.6. Per-key Sequential Consistency
As Section 4.2 pointed out, sequential consistency is non-local, meaning that a protocol that enforces SO and CASL ordering on a per-object basis (termed CASL-per-key) does not necessarily guarantee a global SO and CASL ordering across all operations. In fact, such a protocol implements per-key sequential consistency.
This consistency level was first studied in the PNUTS system (Cooper et al., 2008), a highly-concurrent data serving system that provides per-record consistency. However, modern distributed systems typically have complicated client-side logic layered on top of a non-transactional object store, where each client is interested in more than one object. This makes the object-key-oriented consistency level less appealing than session-oriented causality levels. The photo-album case described in Section 4.3 would be a good example that demonstrates the limitations of per-key sequential consistency.
@@ 5.3. Availability Upper Bounds
Table 2. Availability Upper Bound of Consistency Levels.
Most of these availability bounds have been proven in previous literature (Mahajan et al., 2012; Bailis et al., 2013a). Linearizability, regular sequential consistency, and bounded staleness are obviously weakly available because of the RT constraint or the delay constraint: clients connecting to servers separated on opposite sides of a network partition have no way of knowing the acknowledgment time of operations made on the other side, unless operations on that side are blocked indefinitely. Sequential consistency cannot be sticky available because of its non-locality, as counter-examples similar to the one presented in Section 4.2 can be constructed; in contrast, per-key sequential is sticky available. Bailis et al. have proven that the writes follow reads, monotonic reads, and monotonic writes session guarantees are totally available, while read my writes requires stickiness (Bailis et al., 2013a). Causal and PRAM consistency are therefore both sticky available. Mahajan et al. have proven that real-time causal is as available as causal consistency (given one-way convergence, which is assumed in our model) (Mahajan et al., 2012). Causal+ is also sticky available following this result. Eventual and weak consistency are both totally available: clients can make progress on any live server.

### v4

# A Unified, Practical, and Understandable Model of Non-transactional Consistency Levels in Distributed Replication

@@ Abstract.
We present a practical model of non-transactional consistency levels in the context of distributed data replication. Unlike prior work, our simple Shared Object Pool (SOP) model defines common consistency levels in a unified framework centered around the single concept of ordering. This naturally reflects modern cloud object storage services and is thus easy to understand. We show that a consistency level can be intuitively defined by specifying two types of constraints on the validity of orderings allowed by the level: convergence, which bounds the lineage shape of the ordering, and relationship, which bounds the relative positions between operations. We give examples of representative protocols and systems, and discuss their availability upper bound. To further demonstrate the expressiveness and practical relevance of our model, we use it to implement a Jepsen-integrated consistency checker for the four most common levels (linearizable, sequential, causal+, and eventual); the checker analyzes consistency conformity for small-scale histories of real system runs (etcd, ZooKeeper, and RabbitMQ).
@@ 1. Introduction
The rest of the paper is organized as follows. §2 describes our problem model setup, defines ordering, and explains the meaning of non-transactional consistency within this context. §3 defines all variants of ordering validity constraints. §4 presents the hierarchy of selected consistency levels, dissects their ordering validity guarantees, explains their practical differences, and gives examples of representative protocols and systems. §5 discusses the availability upper bounds in the presence of network partitioning. §6 presents a checker prototype implementation. §7 concludes.
@@ 2.1. Shared Object Pool (SOP) Model
We consider a storage service shared by multiple clients, as shown in Figure 1. The service appears to be a pool of objects. Each object has a unique name and contains a value; it is a register in classic literature. The only way to learn about an object’s value is through the result of a read operation, which we introduce below. Objects are not necessarily stored as physical bytes on physical machines; the SOP model is entirely conceptual and is agnostic to any actual design of protocols and implementation of systems.
Clients are single-threaded, closed-loop entities that invoke operations on the service. When a client $c$ issues an operation $p$, it blocks until the acknowledgment of $p$ by the service. Multi-threaded or asynchronous client implementations should be modeled as multiple SOP clients. An operation is of one of the following types:
@@
The types of objects and operations can be generalized. For example, objects can be counters or queues, and RMW operations can be extended to arbitrary commands. We use the above read-write-style definition throughout this paper for clarity.
The service maintains a possibly partial ordering $O$ of all operations acknowledged. The ordering $O$ captures dependencies between operations enforced by the service and materializes the result of each operation. Given a workload of operations generated by clients, whether an ordering is acceptable or not is decided by some validity constraints. Modeling the validity constraints guaranteed by the service effectively models its interface semantics, hence its consistency level. The following three subsections explain the meaning of workload, ordering, and consistency, respectively.
@@ 2.4. Meaning of Consistency
The consistency level of the storage service is determined by which orderings of operations are considered valid given any physical timeline workload. In other words, the consistency level enforces what validity constraints must be held on the ordering given any workload. A stronger consistency level imposes more constraints than a weaker one and therefore disallows more orderings, exposing an interface that is more restrictive in the protocol design space and in the meantime easier to use by clients. In contrast, a weaker consistency level relaxes certain constraints and opens up new opportunities in the protocol design space, albeit providing weaker semantic guarantees for clients.
An ordering represents logical dependencies among operations, similar to Lamport’s definition of logical clock on events (Lamport, 1978b), and does not necessarily capture physical time; in fact, whether physical time is respected or not is one of the validity constraints that differentiate several consistency levels. Our SOP model shares similarities with the specification framework for replicated data types proposed by Burckhardt et al. (Gotsman et al., 2014); the differences are that we simplify the notion of ordering and cover stronger consistency levels (rather than focusing only on causal and eventual consistency models).
@@ 3.1.2. Convergent Partial Order (CPO)
In addition, the result of a read must be strongly convergent (Viotti and Vukolić, 2016), meaning that it must observe all operations to the same object that immediately precede it. If multiple operations with different values to the same object all immediately precede the read and they are unordered with each other, then the read must return the set of all these values (or a reduced value over the set by applying a convergent reduction function, as described in §2.1).
Below is an example ordering that satisfies CPO (but not SO):
@@ 3.1.3. Non-convergent Partial Order (NPO)
Below is an example ordering that satisfies NPO (but not CPO):
@@ 3.2.2. Causal (CASL)
For instance, the following is an SO ordering that satisfies CASL (but not RT), given the same example timeline of §3.2.1:
@@
Session Guarantees. A popular approach to interpreting causality is to think from each client’s perspective (termed a session (Terry et al., 1994)) and decompose the CASL constraint into four session guarantees:
@@ 3.2.3. First-In-First-Out (FIFO)
For instance, the following is an SO ordering that satisfies FIFO (but not CASL), given the same example timeline of §3.2.1:
@@ 3.2.4. None Relationship
An ordering could, of course, place no restrictions on the relative positions of operations. In this case, operations issued by the same client may get arbitrarily reordered. Writes by the same client may be visible to another client in a different order than issued, and a client’s read may fail to observe its own preceding write.
@@ 4. Consistency Levels
4.1. Linearizability
4.2. Sequential Consistency
For example, given the following physical timeline:
$c:$$d:$$c\text{W}x{\angle}1$$d\text{W}x{\angle}2$$d\text{R}x$
A linearizable ordering must be SO and RT:
$|c\text{W}x{\angle}1|$ $|d\text{W}x{\angle}2|$ $|d\text{R}x{:}2|$
$|d\text{W}x{\angle}2|$ $|c\text{W}x{\angle}1|$ $|d\text{R}x{:}1|$
@@ 4.2. Sequential Consistency
In the example depicted by Figure 3, client $c$ first issues a write of value 1 to object $x$ and waits for its acknowledgment. It then sends a message to client $d$ through an external inter-client channel saying “I have finished my write to $x$ and you can go ahead to operate on $x$.” Client $d$ then issues its own write of value 2 and expects to read out 2 afterwards. However, since the message from $c$ to $d$ is external to the service, a sequentially consistent service may reorder $d$’s write ahead of $c$’s, and return value 1 for $d$’s read.
@@
A service that provides linearizability will be able to capture such implicit external dependencies because of the real-time property, as $|d\text{W}x{\angle}2|$ starts after $|c\text{W}x{\angle}1|$’s acknowledgment in physical time [^6] Note that this is not to be confused with the external consistency property in distributed transaction processing systems (Gifford, 1981; Corbett et al., 2013), which means that transactions are serialized into the same order as their commit order.. In contrast, weaker levels that do not honor real time will at best capture logical causality, the normal definition of causality (CASL) as described in §3.2.2.
@@
Notice that given the result of $d$ reading 1 out of $x$ and $c$ reading 2 out of $y$, it is impossible to resolve an SO and CASL ordering across all six operations. This implies that a protocol that guarantees sequential consistency on each object may fail to come up with a global sequence of operations. In fact, such a protocol provides per-key sequential consistency (see §4.5.6).
@@
Adding support for read-modify-write (RMW) operations to such protocols is a non-trivial task (Burke et al., 2020). In particular, we cannot simply treat RMW operations in the same way as pure writes, because RMWs require a stable base value to determine the result of the read. Systems that demand compare-and-swap (CAS) operations (such as the LogOnce operation on shared logs (Guo et al., 2022)) may have to opt for a service that provides linearizability (or regular sequential consistency (Helt et al., 2021) as discussed in §4.5.1).
@@ 4.4. Eventual Consistency
Eventual consistency, as the name suggests, is a consistency level that only requires reads on an object to return a consistent value if no updates are being made to the object. There is no relationship constraint between operations, meaning that any pair of operations are allowed to get reordered, let alone preserving causality, in the final ordering. Eventual consistency is widely adopted in high-demand systems where high performance, scalability, and availability outweigh the need for timely consistency.
Eventual (Strong) Convergence. Although eventual consistency is sometimes used interchangeably with weak consistency, it does impose one extra requirement on the service: the decided ordering must be (strongly) convergent (Viotti and Vukolić, 2016). Writes eventually become visible to all readers, and reads on the object must all return the same value if they have observed the same writes. This corresponds to the strong eventual variant defined in previous literature (Vogels, 2008); in our model, it is captured by the CPO constraint.
@@
Quiescent Consistency. A closely related, vaguely defined term is quiescent consistency (Herlihy and Shavit, 2011). In a commonly accepted definition, special periods of physical time called quiescence period are identified, during which no write operations are happening. All operations acknowledged ahead of the period are ordered before those that start after the period. Quiescent consistency is weaker than eventual consistency, because it effectively makes no guarantees at all if a system-wide quiescence period never appears (Viotti and Vukolić, 2016).
Protocols & Systems. Eventual consistency is widely adopted by web-scale systems in the form of gossiping protocols and anti-entropy propagation (Schroeder et al., 1984; DeCandia et al., 2007; Elhemali et al., 2022; Lakshman and Malik, 2010). These systems value performance and scalability greatly and can tolerate inconsistencies. A notable line of research related to eventual consistency is on Conflict-free Replicated Data Types (CRDTs) (Letia et al., 2009; Shapiro et al., 2011; Thomson et al., 2012).
@@ 4.5.1. Regular Sequential Consistency
Invariant-equivalence to Linearizability. It is shown that regular sequential consistency is invariant-equivalent to linearizability (Helt et al., 2021), meaning that: 1 it is local (see §4.2) and 2 it inherently supports RMW operations thanks to stable ordering of writes. However, it does not guarantee to capture external causality dependencies, making it still slightly weaker than linearizability. If external causality is not an issue, a linearizable replication system can seamlessly adopt regular sequential consistency to improve the performance of read-only operations.
@@ 4.5.2. Real-time Causal Consistency
Assuming that the system is composed of a set of symmetric message-passing replica nodes, Mahajan et al. have proven in (Mahajan et al., 2012) that real-time causal consistency is the strongest possible level that is achievable in an always-available, one-way convergent system (which is implied by our definition of sticky available in §5).
Fork-based Consistency Models. A family of fork-based consistency models has been developed to deal with Byzantine faults in a system containing untrusted components. For example, a fork-linearizable system ensures that if any two replicas have observed different orderings (i.e., forked by an adversary), then their writes will never be visible to each other afterwards (i.e., they cannot be joined again). Fork causal consistency is a family of consistency levels that weaken causal consistency to tolerate Byzantine replicas and enforce causal consistency among correct replicas (Mahajan et al., 2011).
@@ 4.5.3. Causal Consistency
Causal and causal+ consistency have been explained in §4.3. As a recap, a service that provides causal consistency must give an ordering that is NPO and CASL given any physical timeline workload. Such an ordering captures all the potential causality dependencies between operations, but does not demand convergent conflict resolution, meaning that different clients are allowed to forever retrieve different values from reads on the same object.
As mentioned in §3.2.2, causal consistency can be defined exactly as the combination of the four session guarantees (Brzezinski et al., 2004; Jepsen, 2016).
@@ 4.5.6. Per-key Sequential Consistency
As §4.2 pointed out, sequential consistency is non-local, meaning that a protocol that enforces SO and CASL ordering on a per-object basis (termed CASL-per-key) does not necessarily guarantee a global SO and CASL ordering across all operations. In fact, such a protocol implements per-key sequential consistency.
This consistency level was first studied in the PNUTS system (Cooper et al., 2008), a highly-concurrent data serving system that provides per-record consistency. However, modern distributed systems typically have complicated client-side logic layered on top of a non-transactional object store, where each client is interested in more than one object. This makes the object-key-oriented consistency level less appealing than session-oriented causality levels. The photo-album case described in §4.3 would be a good example that demonstrates the limitations of per-key sequential consistency.
@@ 5.3. Availability Upper Bounds
Table 2. Availability Upper Bounds of Consistency Levels.
Most of these availability bounds have been proven in previous literature (Mahajan et al., 2012; Bailis et al., 2013a). Linearizability, regular sequential consistency, and bounded staleness are obviously weakly available because of the RT constraint or the delay constraint: clients connecting to servers separated on opposite sides of a network partition have no way of knowing the acknowledgment time of operations made on the other side, unless operations on that side are blocked indefinitely. Sequential consistency cannot be sticky available because of its non-locality, as counter-examples similar to the one presented in §4.2 can be constructed; in contrast, per-key sequential is sticky available. Bailis et al. have proven that the writes follow reads, monotonic reads, and monotonic writes session guarantees are totally available, while read my writes requires stickiness (Bailis et al., 2013a). Causal and PRAM consistency are therefore both sticky available. Mahajan et al. have proven that real-time causal is as available as causal consistency (given one-way convergence, which is assumed in our model) (Mahajan et al., 2012). Causal+ is also sticky available following this result. Eventual and weak consistency are both totally available: clients can make progress on any live server.
@@
6. Checker Implementation
To further demonstrate the uniformity, practicality, and understandability of the SOP model, we apply it to consistency checking. Assume a known number of clients using a key-value store service as depicted throughout the paper. Given a history of client operations as input, our consistency model should be able to decide which consistency levels the service conforms to (according to the specific history) and which levels it certainly already violates.
We use Jepsen, a distributed system testing, fault injection, and analysis toolchain (Jepsen, 2025a) (open-sourced by the same-named company (Jepsen, 2016)), as the basis of our implementation. Jepsen offers an automated workflow for running real distributed systems, generating client workloads, injecting failures, recording the execution history, and performing consistency and availability analysis.
We implement a consistency levels conformity checker prototype in ~1k lines of Rust, using SOP orderings as the underlying mechanism. We add ~1k lines of Clojure wrappers to integrate the checker with Jepsen and make it a selectable alternative to the original Knossos linearizability analyzer for key-value operations. The demo can be found at https://github.com/josehu07/jepsen.demo.
6.1. Checker Logic
The checker takes as input from the Jepsen execution stage a history, which is a sequence of events where each event is either the invocation or the completion of a client operation. There are three types of operations: read (R), write (W), and compare-and-swap (CAS). The three types correspond to the model’s definition described in §2.1, with CAS being a concrete, representative type of an RMW operation that conditionally writes a new value if passing an equality check on an expected old value.
The checker outputs four flags to indicate whether the given history conforms to the four common consistency levels: linearizable (Linr.), sequential (Seql.), causal+ (Casl+), and eventual (Evtl.). Conveniently, due to the chain ranking across the four levels, satisfying a higher level guarantees all weaker levels. Note that the result of a run is specific to the particular history produced in the run, and the system in general could be at a weaker level than tested.
The internal logic of the checker goes as follows. 1 It parses the history into a timeline resembling §2.2, stored as a collection of per-client queues of spans, where each span represents a specific operation with start and end timestamps. 2 It repeatedly drains the queues in bulk of concurrent operations, and tries to iterate through all possible constructions of ordering graphs. If a graph satisfying both the convergence and relationship validity constraints of a level is found when all queues in the timeline have been drained, that level is satisfied. The iterative process starts from ”easier” graphs (e.g., SO graphs with RT relationship constraint), seeking stronger levels first to terminate early, before moving on to ”harder” graphs (e.g., CPO graphs with more flexible relationship constraints). 3 If all the CPO graph possibilities are exhausted for all chunks of spans, or if any read returns a corrupted or never-seen value, the checker terminates with all flags set to false (meaning weak consistency).
6.2. Analysis Results
We run the Jepsen workflow on three representative systems: the etcd key-value store (etcd, 2023), the ZooKeeper coordination service (Hunt et al., 2010), and the RabbitMQ message broker (RabbitMQ, 2025), with various configurations when relevant. All systems are run with 5 replicas distributed across 5 CloudLab machines (Duplyakin et al., 2019), and are directed to expose a replicated key-value store service API to clients.
In each run, 10 clients are distributed across the same machines evenly, and generate 30 seconds of workloads with random keys and values at a global 200 ops/sec rate. Network partitioning faults are injected every 10 seconds and last 5 seconds each time. This gives a diverse coverage of the consistency hierarchy across the four most common levels. We do not measure availability.
Table 3 presents the results of the runs. Our SOP-based checker outputs fine-grained consistency validation results that span the four levels. Jepsen’s original Knossos analyzer outputs a binary decision on linearizability only. The results of all six system configurations match what we expected from the system deployments; we explain the deployment modes below.
etcd Modes. etcd is a Raft-consensus-backed, strongly-consistent key-value store for critical data with transaction support. Default deployment uses quorum reads, following the Raft protocol strictly, and is therefore linearizable. If reads are allowed to be acknowledged before reaching a majority quorum (Stale read), they could miss the latest committed writes, and the service degrades to sequential consistency. We also test a mode where CAS operations are implemented manually as serializable transactions instead of single-point operations, which also brings the overall consistency down to sequential.
ZooKeeper (ZK) Modes. ZooKeeper is a sequentially-consistent coordination service backed by the ZAB primary-backup protocol. We use ZK through the Avout library, which provides a distributed Clojure atom abstraction using ZK as access locks. Although sequentially consistent, triggering a non-linearizable read result is rare (as it requires stale locks to be held in close succession); thus, our test run yielded a linearizable history (). We also include a mode where a subset of atoms is replaced with local atom references without cross-node communication. This caps those atoms at causal+ consistency.
RabbitMQ (RMQ) Modes. RabbitMQ is a message queueing and brokerage system. We build a peer-to-peer broadcasting layer using RabbitMQ queues co-located with each node as the communication media between them. This resembles a weakly-consistent key-value store service to clients, where updates are propagated lazily to peers through background announcements.
Performance Limitation. The main purpose of the checker implementation is to demonstrate the uniformity, expressiveness, and practical relevance of the SOP model. A major limitation lies in brute-force ordering graph construction, which has sub-factorial complexity w.r.t. the number of concurrent writes (bounded by the number of clients), leading to long analysis time and high memory consumption. Checking for linearizability alone takes ~1s, on par with Jepsen’s in our small-scale tests, but weaker levels require hours to explore. In practice, level-specific algorithms are used (Horn and Kroening, 2015; Jepsen, 2025b; Athalye, 2017; Papadimitriou, 1979; Bouajjani et al., 2017; Adya, 1999) with auxiliary information from the tested system (such as versions and the believed serialization order).
System Configuration SOP-based Checker Jepsen
System Mode Linr. Seql. Casl+ Evtl. Knossos
etcd Quorum read Pass
Stale read No
CAS as txns No
ZK Locked atoms Pass
Local refs No
RMQ P2P announce No
Table 3. Jepsen Workflow Checker Analysis Results.